# Strong primeness for equivalence relations arising from Zariski dense subgroups

### v1

@@ 3.1. Zimmer’s method for product equivalence relations
The reasoning in the first paragraph shows that for every $u\in U$, there exist $F\in\mathscr{G}$ and $b\in G$ such that $(\varphi(u,v)H_{u},\psi(u,v)J_{u})=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$. This means exactly that $(u,b,F)\in B$. Therefore, we have $\pi_{U}(B)=U$. Using (2.7), it follows that $B\subset U\times G\times\mathscr{G}$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $U\to G\times\mathscr{G}:u\mapsto(b_{u},F_{u})$ such that $(u,b_{u},F_{u})\in B$ for $\nu$-almost every $u\in U$. Thus, we may choose Borel maps $\theta:W\to G$ and $\mathfrak{L}:U\mapsto\operatorname{Sub}(G):u\mapsto L_{u}$ such that $\theta(u,v)=F_{u}(v)$ for $\zeta$-almost every $(u,v)\in W$ and $L_{u}=H_{u}\cap b_{u}J_{u}b_{u}^{-1}$ for $\nu$-almost every $u\in U$. Note that for $\nu$-almost every $u\in U$, letting $\mathbf{L}_{u}=\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1}<\mathbf{G}$, which is a $k$-subgroup, we have $L_{u}=\mathbf{L}_{u}(k)$. For $\nu$-almost every $u\in U$, using the $G$-equivariant homeomorphism
@@ 3.2. Proof of Theorem A
The proof of Claim 3.5 consists of two steps. Firstly, we consider the product equivalence $\mathscr{T}_{1}\times\mathscr{U}_{2}$ together with the subequivalence relation $\mathscr{U}_{1}\times\mathscr{U}_{2}\leq\mathscr{T}_{1}\times\mathscr{U}_{2}$ and the $(\mathscr{U}_{1}\times\mathscr{U}_{2},\alpha|_{\mathscr{U}_{1}\times\mathscr{U}_{2}})$-equivariant measurable map $f:Z\to G/H$. By Theorem 3.3$(\rm ii)$, we may choose a minimal pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ for which there exists a Zariski connected $k$-subgroup $\mathbf{J}<\mathbf{G}$ such that $\mathbf{J}=\mathbf{J}_{z_{1}}^{0}$ for $\zeta_{1}$-almost every $z_{1}\in Z_{1}$. By Claim 3.4, there exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and a $(\mathscr{U},\alpha|_{\mathscr{U}})$-equivariant measurable map $f:Z\to G/H$. Choose a Borel section $\sigma:G/H\to G$. Consider the measurable map $\varphi=\sigma\circ f:Z\to G$ and the constant map $\mathfrak{H}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto H$. Then $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$. By minimality of $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ and since $\mathbf{H}<\mathbf{G}$ is a proper $k$-subgroup, it follows that $\mathbf{J}<\mathbf{G}$ is a proper $k$-subgroup.
@@
satisfies $\pi_{W}(B)=W$. Using (2.6), it follows that $B\subset W\times G$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ such that $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for $\zeta_{1}$-almost every $z_{1}\in W$. Upon discarding a null Borel subset, we may assume that the Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ satisfies $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for every $z_{1}\in W$. We may extend the Borel map $Z_{1}\to G:z_{1}\mapsto b_{z_{1}}$ by declaring that $b_{z_{1}}=e$ for every $z_{1}\in Z_{1}\setminus W$. Upon replacing the pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ by the pair $(\rho,\mathfrak{K})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$, where $\rho:Z\to G:(z_{1},z_{2})\mapsto\psi(z_{1},z_{2})b_{z_{1}}^{-1}$ and $\mathfrak{K}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}$, we may assume that $J_{z_{1}}<K$ for every $z_{1}\in W$.
@@ 3.5. Proof of Theorem D
It remains to prove that $T_{1}\cap T_{2}=\emptyset$. By contradiction, assume that $T_{1}\cap T_{2}\neq\emptyset$ and let $i\in T_{1}\cap T_{2}$. Set $\Lambda=\prod_{i\neq j}\Gamma_{i}$ and $\mathscr{S}=\mathscr{R}(\Lambda\curvearrowright X)$. Arguing as in the proof of Corollary B, by minimality of $T_{1}$ and $T_{2}$ and [DHI16, Lemma 2.8(2)], we have that $P_{j}\npreceq_{M}^{s}\operatorname{L}(\Lambda\curvearrowright X)$ for every $j\in\{1,2\}$. By Fact 2.2$(\rm i)$, this further implies that $\mathscr{T}_{j}\npreceq_{\mathscr{R}}\mathscr{S}$ as subequivalence relations. This however contradicts Theorem 3.6. Therefore, we have $\{1,\dots,n\}=T_{1}\sqcup T_{2}$. ∎

### v2

@@ 3.1. Zimmer’s method for product equivalence relations
The reasoning in the previous paragraph shows that for every $u\in U$, there exist $F\in\mathscr{G}$ and $b\in G$ such that $(\varphi(u,v)H_{u},\psi(u,v)J_{u})=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$. This means exactly that $(u,b,F)\in B$. Therefore, we have $\pi_{U}(B)=U$. Using (2.7), we have that $B\subset U\times G\times\mathscr{G}$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $U\to G\times\mathscr{G}:u\mapsto(b_{u},F_{u})$ such that $(u,b_{u},F_{u})\in B$ for $\nu$-almost every $u\in U$. Thus, we may choose Borel maps $\theta:W\to G$ and $\mathfrak{L}:U\mapsto\operatorname{Sub}(G):u\mapsto L_{u}$ such that $\theta(u,v)=F_{u}(v)$ for $\zeta$-almost every $(u,v)\in W$ and $L_{u}=H_{u}\cap b_{u}J_{u}b_{u}^{-1}$ for $\nu$-almost every $u\in U$. Note that for $\nu$-almost every $u\in U$, letting $\mathbf{L}_{u}=\mathbf{H}_{u}\cap b_{u}\mathbf{J}_{u}b_{u}^{-1}<\mathbf{G}$, which is a $k$-subgroup, we have $L_{u}=\mathbf{L}_{u}(k)$. For $\nu$-almost every $u\in U$, using the $G$-equivariant homeomorphism
@@ 3.2. Proof of Theorem A
The proof of Claim 3.5 consists of two steps. Firstly, we consider the product equivalence $\mathscr{T}_{1}\times\mathscr{U}_{2}$ together with the subequivalence relation $\mathscr{U}_{1}\times\mathscr{U}_{2}\leq\mathscr{T}_{1}\times\mathscr{U}_{2}$ and the $(\mathscr{U}_{1}\times\mathscr{U}_{2},\alpha|_{\mathscr{U}_{1}\times\mathscr{U}_{2}})$-equivariant measurable map $f:Z\to G/H$. By Theorem 3.3$(\rm ii)$, we may choose a minimal pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ for which there exists a Zariski connected $k$-subgroup $\mathbf{J}<\mathbf{G}$ such that $\mathbf{J}=\mathbf{J}_{z_{1}}^{0}$ for $\zeta_{1}$-almost every $z_{1}\in Z_{1}$. By Claim 3.4, there exist a proper $k$-subgroup $\mathbf{H}<\mathbf{G}$ and a $(\mathscr{U},\alpha|_{\mathscr{U}})$-equivariant measurable map $f:Z\to G/H$. Choose a Borel section $\sigma:G/H\to G$. Consider the measurable map $\varphi=\sigma\circ f:Z\to G$ and the constant map $\mathfrak{H}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto H$. Then for $\zeta_{1}$-almost every $z_{1}\in Z_{1}$, the Borel map $Z_{2}\to G/H:z_{2}\mapsto\varphi(z_{1},z_{2})H=f(z_{1},z_{2})$ is $(\mathscr{U}_{2},\alpha_{z_{1}})$-equivariant and so we may regard $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$. By minimality of $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ and since $\mathbf{H}<\mathbf{G}$ is a proper $k$-subgroup, it follows that $\mathbf{J}<\mathbf{G}$ is a proper $k$-subgroup.
@@
satisfies $\pi_{W}(B)=W$. Using (2.6), we have that $B\subset W\times G$ is a Borel subset. By the measurable selection theorem (see Theorem 3.1), there exists a Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ such that $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for $\zeta_{1}$-almost every $z_{1}\in W$. Upon discarding a null Borel subset, we may assume that the Borel map $W\to G:z_{1}\mapsto b_{z_{1}}$ satisfies $b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}<K$ for every $z_{1}\in W$. We may extend the Borel map $Z_{1}\to G:z_{1}\mapsto b_{z_{1}}$ by declaring that $b_{z_{1}}=e$ for every $z_{1}\in Z_{1}\setminus W$. Upon replacing the pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$ by the pair $(\rho,\mathfrak{K})\in\mathfrak{A}(\alpha|_{\mathscr{T}_{1}\times\mathscr{U}_{2}})$, where $\rho:Z\to G:(z_{1},z_{2})\mapsto\psi(z_{1},z_{2})b_{z_{1}}^{-1}$ and $\mathfrak{K}:Z_{1}\to\operatorname{Sub}(G):z_{1}\mapsto b_{z_{1}}J_{z_{1}}b_{z_{1}}^{-1}$, we may assume that $J_{z_{1}}<K$ for every $z_{1}\in W$.
@@ 3.5. Proof of Theorem D
It remains to prove that $T_{1}\cap T_{2}=\emptyset$. By contradiction, assume that $T_{1}\cap T_{2}\neq\emptyset$ and let $i\in T_{1}\cap T_{2}$. Set $\Lambda=\prod_{j\neq i}\Gamma_{j}$ and $\mathscr{S}=\mathscr{R}(\Lambda\curvearrowright X)$. Arguing as in the proof of Corollary B, by minimality of $T_{1}$ and $T_{2}$ and [DHI16, Lemma 2.8(2)], we have that $P_{j}\npreceq_{M}^{s}\operatorname{L}(\Lambda\curvearrowright X)$ for every $j\in\{1,2\}$. By Fact 2.2$(\rm i)$, this further implies that $\mathscr{T}_{j}\npreceq_{\mathscr{R}}\mathscr{S}$ as subequivalence relations. This however contradicts Theorem 3.6. Therefore, we have $\{1,\dots,n\}=T_{1}\sqcup T_{2}$. ∎